Eliminate the arbitrary function and hence obtain the partial differential equation:

φ (
z
x
3
,
y
x
) = 0
1
Expert's answer
2021-01-14T17:32:34-0500

"\phi=(zx^3,yx)=0"

"zx^3=c_1"

"yx=c_2"


"dz=-\frac{3c_1dx}{x^4}, dx=-\frac{c_2dy}{y^2}\implies dz=\frac{3c_1c_2dy}{x^4y^2}"


"dz=-\frac{3c_1dx}{x^4}=\frac{3c_1c_2dy}{x^4y^2}"


"\frac{dz}{3c_1c_2}=-\frac{dx}{c_2x^4}=\frac{dy}{x^4y^2}"


"x^4y^2q-c_2x^4p=3c_1c_2"